factor the quadratic expression completely, and find the roots of the expression.
20x^2 + 13x + 2

Answer by checkley71(8403)

20X~2+13X+2 OR (5x+2)(4X+1)=0 OR 5X+2=0 OR 5X=-2 OR X=-2/5 & 4X+1=0 OR 4X=-1 OR X=-1/4